In 7 years, maria will be 4 times as old as angelica will be then. The sum of their ages now is 71. How old will each of them be in 5 years?
Answer by stanbon(75887)

In 7 years, maria will be 4 times as old as angelica will be then. The sum of their ages now is 71. How old will each of them be in 5 years?
------------------------------
In-Seven-Years DATA:
Angelica's Age: x yrs old
Maria's Age: 4x yrs old
-------------------------------
Their Ages NOW DATA:
Angelica: x-7 yrs old
Maria's Age: 4x-7 yrs old
-------------------------------
Equation:
The sum of their ages now is 71
x-7 + 4x-7 = 71
5x - 14 = 71
5x = 85
x = 17
x-7 = 10 yrs old (Angelica's age NOW
4x-7= 68 yrs old (Maria's age NOW
--------------------
Their Ages 5 years from Now:
Angelica will be 10+5 = 15 yrs old
Maria will be 68+5 = 73 yrs old
=======================================
Cheers,
Stan H.